OpenStudy (anonymous):

what does it mean when a problem says "in terms of" example Express the following logarithm as specified . Natural log Root 40.5 in terms of natural log 3 and natural log 4. Thank you in advance for your help.

OpenStudy (anonymous):

By natural log 3, do you mean \(\log_{e}(3)=\ln(3)\) ?

OpenStudy (jdoe0001):

in terms of, meaning the answer will be exclusively expressed in those tems

OpenStudy (jdoe0001):

thus ... far all I've found is ... \(\bf ln(\sqrt{40.5}) \implies ln\left(40.5^{\frac{1}{2}}\right) \implies \cfrac{1}{2}ln(40.5) \implies \cfrac{ln(40.5)}{2}\\\qquad \\ \textit{now we can say that } 40.5 = 3^2 \times 4.5 \textit{thus}\\ \qquad \\ \cfrac{ln(40.5)}{2} \implies \cfrac{ln(3^2 \times 4.5)}{2} \implies \cfrac{ln(3^2) + ln(4.5)}{2}\\ \implies \cfrac{2ln(3) + ln(4.5)}{2} \implies ln(3) + \cfrac{ln(4.5)}{2}\)

OpenStudy (anonymous):

For some reason my computer is having difficulties in rendering LaTex just give me a moment.

OpenStudy (jdoe0001):

@KeithAfasCalcLover try maybe at http://mathb.in/ it has better latency

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."